The energy of light calculated using the formulas
Answer #1

E =h*c/\lambda

= 6.26*10^-34 J.s * 3.*10^8 m/s/5.6*10^2 *10^-9 m

= 3.35*10^-19 J = 3.35*10^-16 kJ

energy of 1 mol of photon = 3.35*10^-16 *6.023*10^23 = 20.18 *10^7 kJ/mol

----------------------------------------------------

E = 266kJ/mol

energy of one photon = 266*1000 J/6.023*10^23 = 44.16*10^-20 J/mol

wavelength = 6.26*10^-34 J.s * 3.*10^8 m/s/ 44.16*10^-20 J/mol

= 0.425 *10^-6 m = 425 nm